OpenStudy (anonymous):

Is the set of whole numbers the same as the set of positive integers? Explain.

OpenStudy (unklerhaukus):

it depends on your definitions, is zero in the set?

OpenStudy (anonymous):

A number without fractions is a whole number: 0,1,2,3,4,5....... Positive integers are all the whole numbers greater than zero: 1, 2, 3, 4, 5,...

OpenStudy (anonymous):

@UnkleRhaukus these are my definitions : A. No; the set of positive integers includes 0 but the set of whole numbers does not. B. Yes; both sets start at 1 and continue into infinity. C. No; the set of whole numbers includes 0 but the set of positive integers does not. D. Yes; both sets start at 0 and continue into infinity.

OpenStudy (anonymous):

whole numbers are 1,2,3....... integers are 0,1,2,3........

OpenStudy (unklerhaukus):

some definition of whole numbers includes zero some definitions do not

OpenStudy (anonymous):

@UnkleRhaukus i dont think whole numbers includes zero in some definitions

OpenStudy (anonymous):

are u here @UnkleRhaukus

OpenStudy (unklerhaukus):

here?

OpenStudy (anonymous):

A set of whole number always includes zero

OpenStudy (dape):

The integers is the set \[ \mathbb{Z} = \{...,-2,-1,0,1,2,...\} \]

OpenStudy (anonymous):

If zero is removed then it becomes set of natural numbers

OpenStudy (anonymous):

@UnkleRhaukus and @uzumakhi

Parth (parthkohli):

Here's my reasoning: 1) 0 is not in the set of positive integers. 2) Why would anyone bother to create two different sets with the same elements? Think about it!

OpenStudy (dape):

No, the naturals are the set \[ \mathbb{N} = \{0,1,2,...\} \]

mathslover (mathslover):

0 is \(neither\) positive nor negative..

OpenStudy (anonymous):

C. No; the set of whole numbers includes 0 but the set of positive integers does not.

OpenStudy (anonymous):

^ thats the correct answer

OpenStudy (anonymous):

@sauravshakya zero is present in natural numbers

mathslover (mathslover):

@uzumakhi *no*

OpenStudy (anonymous):

No. @uzumakhi

Parth (parthkohli):

@uzumakhi Nuh-uh.

OpenStudy (dape):

The positive integers are \[ Z^+=\{1,2,...\} \] The whole number is simply the same as the integers, which includes the negatives and zero.

mathslover (mathslover):

Zero is present in whole numbers. --> correction

Parth (parthkohli):

Looks like you gotta read more set theory. Eh? @dape That's right.

Parth (parthkohli):

"+" signifies the absence of zero.\[\mathbb{R}^{+} \implies \text{Set of reals excluding zero.}\]

Parth (parthkohli):

So, \(\mathbb{R}^{+}\) is the solution in,\[{0 \over x} = 0\]We may say that \(x \in \mathbb{R}^{+}\)

Parth (parthkohli):

Wow. I just started a lecture on Set Theory. lol

OpenStudy (anonymous):

i think C is correct

OpenStudy (dape):

Usually "*" specifies the absence of zero (or more generally, the origin), and "+" specifies the part greater than zero, so that for example \[ \mathbb{R}^+=\{x\in\mathbb{R}|x>0\} \] and \[ \mathbb{C}^*=\mathbb{C}-\{0\}\]

OpenStudy (unklerhaukus):

hmm

OpenStudy (anonymous):

@uzumakhi So u got it?